For the following three vectors, what is 3.С. . (3.4 × B) ? I am messing something up as my answer is not being accepted, and I believe it has to do with my signs. My work with the given values is posted below:

nea 38) Given : A一2.001+4.00 3:00 k B=-A00个43.001+3.00k -6.oot-7.003.-
Answer #1

You are messing up in the matrix calculation signs.

A = 2i + 4j - 3k

B = -2i + 3j + 3k

C = 6i - 7j

3A = 6i + 12j - 9k

3A\times B = \begin{vmatrix} i & j &k \\ 6 & 12 & -9\\ -2 & 3 & 3 \end{vmatrix}

3A\times B = i[(12)(3) - (3)(-9)] - j[(6)(3) - (-2)(-9)] + k[(6)(3) - (-2)(12)]

3A\times B = i[36 - (-27)] - j[18 - 18] + k[18 - (-24)]

3A\times B = 63i + 0j + 42k

3C = 18i - 21j

3C(3A\times B) = (18\times 63)(i.i) + (-21\times 0)(j.j) + (0\times 42)(k.k)

3C(3A\times B) = 1134